3. In the country of Pristine, the labour demand and supply are given by L d = 30 − 4 W and L s = −6 + 2 W correspondingly. Suppose that ancient labour laws dictate that the minimum wage in Pristine is $5. What is the impact of this minimum wage law on Pristine’s labour market?
Answer #1

At wage of $5,

Labor demanded = 30-4(5) = 10 units

Labor supplied = -6 + 2(5) = 4 units

Since at this wage, labor demanded exceeds labor supplied, it leads to a situation of shortage of labor in the market, which will put pressure on wages to increase.